question: Cookie Monster always consumes cookies and milk in fixed proportions: 1 cookies to 3 ounce of milk. If he has an income of $100 and the price per cookie is $1 and the price of an ounce of milk is $3, what is the optimal bundle that Cookie Monster should consume?
Answer #1

We know Cookies=1/3ounces of milk I.e. 3C=Milk

And budget P1X1+ P2X2=M

C+3M=100

10C=100 C=10 and milk=30

Thus Cookie=10 and Ounces of milk=10*3=30

Thus 10 cookies and 30 ounces of milk should be consumed